OpenStudy (narissa):

Please check. What is the greatest common factor?10x^4y^5+15x^3y^6 My answer-->6 is the greatest common factor.

OpenStudy (narissa):

@AbdullahM

OpenStudy (abdullahm):

\(\Large\bf 10x^4y^5\) \(\Large\bf 15x^3y^6\) First factor each one. Do you know how to do that?

OpenStudy (narissa):

No :/

OpenStudy (abdullahm):

For example: \(\Large\bf 10x^4y^5 = 2\cdot 5\cdot x\cdot x\cdot x\cdot x \cdot y\cdot y \cdot y \cdot y \cdot y\) Can you do it for the other one?

OpenStudy (narissa):

3*5*x*x*x*y*y*y*y*y*y

OpenStudy (narissa):

@AbdullahM

OpenStudy (abdullahm):

Now you can see that both share a 5. We can factor that out. How many x's do they both share? For example, how many can you take out maximum from one of them that allows you take out the same amount from the other? How many y'x do they both share?

OpenStudy (abdullahm):

y's**

OpenStudy (narissa):

7x's 11y's

OpenStudy (narissa):

is the greatest common factor 1? because 1 is the greatest number that divides evenly into all of them.

OpenStudy (abdullahm):

No, look at this:  We can take out 3 x's and 5 y's Do you understand?